The magnitude of the acceleration of an object moving along a straight line is equal to the?

A. slope of the displacement time graph

B. Area below the force-time graph

C. Slope of its velocity-time graph

D. Area under its velocity-time graph


1
Expert's answer
2021-11-01T17:13:31-0400

Acceleration ("\vec a") is the derivative of velocity ("\vec v") with respect to time (t).

i.e. "\vec a = \frac {d(\vec v)} {dt}" (its unit is: "ms^{-2}")


That is, "\frac {velocity \ (\vec v)} {time \ (t)} = \frac {ms^{-1}} {s} = ms^{-1-1} = ms^{-2}" = acceleration "(\vec a)"


Therefore, the magnitude of the acceleration "|\vec a|" of an object moving along a straight line is the slope/gradient ("m = \frac {d} {dt}") of its velocity-time graph.

The answer is C.